OpenStudy (anonymous):

What is 120% of 31? Round to the nearest whole number if necessary.

OpenStudy (anonymous):

give me a medal i will solve

OpenStudy (anonymous):

is 37 an option?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

120% from 31=120/100*31=6*31/10

OpenStudy (anonymous):

That's the answer

OpenStudy (anonymous):

120% from 31=120/100*31=6*31/5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

37.2 exact

OpenStudy (anonymous):

it says 'round to the nearest whole number'...

OpenStudy (anonymous):

i know dear